OpenStudy (anonymous):

-0.3(x-4)+x=0.5(3-x)

OpenStudy (jwheele1):

-0.3(x-4)+x=0.5(3-x) distribute both sets -.3x -.12 +x = -2.5 -.5x combine like terms .7x -.12 = -2.5 -.5x isolate variable on one side by adding.5x to both sides 1.2x - .12 = -2.5 isolate variable completely by adding .12 to both sides 1.2x = -2.38 divide both sides by 1.2 x = -1.9833 (33 goes on forever)

OpenStudy (jwheele1):

plug answer in for x to verify that both sides equal the same number in the end...which is called a balance statement

OpenStudy (anonymous):

-0.3x + 1.2 +x =1.5 -0.5x 0.5x -0.3x + x = 1.5 -1.2 1.2x = 0.3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

is im right?